# The Basis of Foot-Sortable Sock Orderings

### v1

@@ Abstract.
A sock ordering is a finite string of elements, each of which has a color. Foot-sorting is a non-deterministic sorting algorithm with the aim of grouping all socks of the same color together, under the constraints of a stack structure; a sock ordering is foot-sortable if there is some way to apply the foot-sorting algorithm such that this aim is realizable. In introducing sock orderings and the foot-sorting algorithm, Defant and Kravitz ask whether there is some list of patterns which all foot-sortable sock orderings must avoid, thus characterizing which sock orderings are foot-sortable. We enumerate this list, the basis of foot-sortable sock orderings, and show that this basis is infinite. We also describe a deterministic polynomial-time algorithm for sorting socks.
@@ 1. Introduction
The idea of sorting permutations, introduced by Donald Knuth in [2], considers sorting a given permutation of $\{1,2,\dots,n\}$ to a desired output under the constraints of a stack structure. In [2], Knuth showed that a permutation is sortable if and only if the permutation avoids the pattern 231. In other words, the permutation $a_{1}a_{2}\cdots a_{n}$ is sortable using the stack structure if and only if there is no set of indices $1\leq i<j<k\leq n$ such that $a_{k}<a_{i}<a_{j}$.
Extending Knuth’s findings, Defant and Kravitz introduced the concept of foot sorting for particular patterns in [1]. This alteration allows for single orderings, called sock orderings, to contain multiple socks of the same color, whereas the original work dealt with orderings containing the elements $\{1,2,\dots n\}$ exactly once. A significant difference in sorting a sock ordering when compared to a permutation is that in sock orderings, there is no pre-specified sorted word which the foot-sorting algorithm must produce. Because of the freedom in output, sorting a sock ordering that is given to be foot-sortable may be easier than sorting a permutation pattern.
Defant and Kravitz prove a similar result to Knuth in [1]: that a sock ordering $\rho$ is foot-sortable if and only if there is some word $\nu$ representing $\rho$ such that $\nu$ avoids the pattern 231. While for sock orderings with only a few colors, one could search for such a word which avoids 231, the process of generating all possible words and checking whether they avoid 231 is very computationally intensive. Allowing for multiple iterations of the stack-sorting procedure, Xia investigated a deterministic foot-sorting algorithm in [3] which hinges on pattern avoidance within the stack, though our work focuses on the sock orderings which are foot-sortable in a single iteration of the sorting procedure. Recently, [4] provided a polynomial-time algorithm for deciding the foot-sortability of an arbitrary sock ordering; we later offer a deterministic algorithm which operates in polynomial time and classifies sock orderings as sortable or unsortable, successfully sorting the orderings for which it is possible.
Defant and Kravitz introduce the idea of a basis of foot-sortable sock orderings in [1]. The basis of foot-sortable sock orderings is the smallest set of sock orderings such that a sock ordering $X$ is foot-sortable if and only if no member of the set is a subpattern of $X$; we introduce several key ideas in relation to this basis and subpatterns in the following section. Defant and Kravitz pose two interesting questions about the basis of foot-sortable sock orderings in [1]:
@@
    Is it possible to enumerate the basis of foot-sortable sock orderings?
@@
    Is the basis of foot-sortable sock orderings finite?
In Section 2, we will define a infinite set, $\Gamma$, of minimally unsortable sock orderings. The primary result of our work is the following theorem which proves that $\Gamma$ is the basis for foot-sortable sock orderings.
Theorem 1.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
The proof of Theorem 1 leads naturally to a deterministic algorithm which we present in Section 3.
2. Notation and Preliminaries
Before enumerating the basis of foot-sortable sock orderings, we lay out some important definitions regarding sock orderings and the patterns they contain as well as provide some examples of pattern containment.
@@ 2. Notation and Preliminaries
A sock ordering $X$ is a finite sequence of socks, each of which has a specified color.
We will use $|X|$ to denote the length of a sock ordering $X$, the number of socks in $X$. Throughout our work, we will use capital letters to denote sock orderings, while we use lowercase letters to denote specific socks or colors. At times, we refer to sock orderings as strings in relation to our algorithms.
Definition 3.
Let $X=x_{1}x_{2}\cdots x_{n}$ and $Z=z_{1}z_{2}\cdots z_{n}$ be sock orderings (so $|X|=|Z|$), and let $\mathcal{A}_{X}$ and $\mathcal{A}_{Z}$ be the sets of distinct sock colors appearing in $X$ and $Z$ respectively. $X$ is equivalent to $Z$ (denoted $X\cong Z$) if there exists a bijective mapping $\varphi:\mathcal{A}_{X}\rightarrow\mathcal{A}_{Z}$ such that $\varphi(x_{i})=z_{i}$ for all $1\leq i\leq n$.
Definition 4.
Let $X$ and $Y$ be sock orderings. We say that $X$ contains $Y$ and write $Y\subseteq X$ if deleting zero or more socks from $X$ yields exactly $Y$, and we say $a\in X$ if $a\subseteq X$. If $Y\subseteq X$, we say that $Y$ is a subordering of $X$. In contrast, we say that $Y$ is a subpattern of $X$ and write $Y\preceq X$ if there exists a subordering $Z\subseteq X$ of $X$ such that $Z\cong Y$. If $Y$ is not a subpattern of $X$, then we write $Y\npreceq X$ and say that $X$ avoids $Y$. For a color $a\in X$, we let $X-a$ be the subordering of $X$ formed by removing every sock of color $a$ from $X$.
Because of the many equivalent ways to represent one sock ordering, it may not always be immediately clear whether a sock ordering is a subpatterns of another. The following example will illustrate:
Example 5.
$babc\preceq abacbab$.
We have $babc\nsubseteq abacbab$, but $abac\subseteq abacbab$ and $abac\cong babc$. Therefore, $babc\preceq abacbab$.
To sort, or attempt to sort, sock orderings, we utilize the foot-sorting algorithm proposed by [1], and we outline the rules here.
Definition 6.
foot-sorting is a non-deterministic algorithm which operates on a sock ordering $X$. At any point in the foot-sorting process, one may choose to:
  • Stack a sock: Move the leftmost sock of the remaining input ordering to the top of the stack, or
  • Unstack a sock: Remove the outermost sock on the foot (the topmost element of the stack) and place it on the right of the output.
Though we utilize the idea of foot-sorting and socks, throughout, we will refer to the sorting process using primarily stack-centered language.
At some points, this algorithm has set steps. For example, the first step in foot-sorting always must be to place the first sock, if it exists, into the stack. Similarly, once nothing remains of the original input, all socks in the stack must be removed beginning with the topmost element of the stack and moving downward.
We denote a particular stage in the foot-sorting process as $(S,R)$. $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot, in the language of [1]), while $R$ represents what remains of the input ordering. Before any sorting steps are applied, we may say that an input ordering $X$ appears as $(\emptyset,X)$. To illustrate this process and notation, we provide an example of foot-sorting here.
Example 7.
Using the foot-sorting algorithm, we may obtain the output $aaccbb$ from the input sock ordering $abcabc$.
  1. (1)
    We equivalently write $abcabc$ as $(\emptyset,abcabc)$. Move the leftmost sock of the remaining input ordering, $a$, to the top of the stack. This yields $(a,bcabc)$.
  2. (2)
    Remove the topmost sock of the stack, $a$, and place it on the right of the output, which is currently empty. We now have $(\emptyset,bcabc)$ with output $\kappa=a$.
  3. (3)
    Move the leftmost sock of the remaining input ordering, $b$, to the top of the stack, leaving $(b,cabc)$ and output $\kappa=a$.
  4. (4)
    Stack $c$ to leave $(bc,abc)$ and output $\kappa=a$.
  5. (5)
    Stack $a$ to leave $(bca,bc)$ and output $\kappa=a$.
  6. (6)
    Unstack $a$ to leave $(bc,bc)$ and output $\kappa=aa$.
  7. (7)
    Unstack $c$ to leave $(b,bc)$ and output $\kappa=aac$.
  8. (8)
    Stack $b,c$, leaving $(bbc,\emptyset)$ and output $\kappa=aac$.
  9. (9)
    Unstack, in order, all remaining socks from the stack to leave (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) and output $\kappa=aaccbb$.
Definition 8.
A stage of the sorting algorithm $(S,R)$ is foot-sortable (or, when clear, sortable) if there is some way to apply the foot-sorting algorithm on $(S,R)$ such that, in the output of the foot-sorting process, all socks of the same color occur consecutively. That is, for some $(S,R)$ containing only the colors $\{a_{1},a_{2},\dots a_{n}\}$, $(S,R)$ may be sorted to $\kappa=a_{\pi(1)}\cdots a_{\pi(1)}a_{\pi(2)}\cdots a_{\pi(2)}\cdots a_{\pi(n)}\cdots a_{\pi(n)}$ where $\pi$ is a permutation of $\{1,\dotsc,n\}$. If there is no such way to apply the foot-sorting algorithm, $X$ is unsortable. We say that a sock ordering $X$ is foot-sortable (resp. unsortable) if $(\emptyset,X)$ is foot-sortable (resp. unsortable).
We see that $abcabc$ in Example 7 is a sortable sock ordering.
The goal of foot-sorting is to begin with an input sock ordering $X$ and apply the steps of the algorithm so that nothing remains in the stack or on the right side. While it can be difficult to decide whether a particular sock ordering is sortable or unsortable, there is an important indicator which can reveal if applying the foot-sorting algorithm has already failed. Note that a sortable sock ordering $X$ may be sorted to a word $\kappa$ such that $aba\not\preceq\kappa$. Due to the nature of the stack structure, once a sock, $a$, appears in the stack, $a$ must appear in the output $\kappa$ before any socks appearing below $a$ in the stack and after any socks appearing above $a$ in the stack. For $(S,R)$ appearing at some point while applying the foot-sorting algorithm, let $S=s_{1}s_{2}\cdots s_{n}$ and set $S^{-1}=s_{n}\cdots s_{2}s_{1}$, the reverse of $S$. It must be that $S^{-1}\subseteq\kappa$. Since in a sorted word $\kappa$, we have $aba\not\preceq\kappa$, and $(aba)^{-1}=aba$, then if $aba\preceq S$, it is not possible to successfully sort the sock ordering. Because of this, if we ever have $aba\preceq S$ while attempting to sort a particular sock ordering, we will immediately stop the sorting attempt as it is not possible to finish sorting after this point. We refer to this $aba$ pattern of distinct colors $a,b$ as a sandwich.
Sandwiches, the colors they contain, and where they appear in sock orderings arise frequently in our work. We may say that the sandwich $aba$ blocks, precedes, or appears before the color $c$, for distinct colors $a,b,c$, at some stage $(S,R)$ if $abac\subseteq SR$. Additionally, we are sometimes concerned about the first sandwich $aba$ which blocks a color $c$. If two distinct sandwiches, $aba$ and $xyx$, precede $c$, then if $a$ and $x$ are distinct colors, $aba$ appears earlier than $xyx$ if the second denoted $a$ appears before the second denoted $x$; if $x$ is $a$, then $aba$ appears earlier than $aya$ if the denoted $b$ appears before the denoted $y$. For example, in $abcabd$, all of the sandwiches $aba,aca,bcb,bab$ block $d$, but $aba$ is the first sandwich which blocks $d$. Finally, we may say $x$-sandwich to describe a sandwich which contains a sock of color $x$.
Interestingly, though it is necessary to avoid $aba$ appearing in the stack while applying the foot-sorting algorithm, a deterministic algorithm which prevents the appearance of sandwiches in the stack does not succeed in sorting all sortable sock orderings; Xia investigated such an algorithm, $\phi_{aba}$, in [3]. We later introduce a deterministic algorithm $\phi_{G}$ which, as a consequence of its operation, prevents $aba$ from appearing in the stack, but has further restrictions that differentiate it from $\phi_{aba}$. It was also shown that $\phi_{aba}$ in [3] may sometimes sort sock orderings which have subpatterns that are unsortable under $\phi_{aba}$. In foot-sorting, however, this is not the case; if there exists no set of steps which may sort some subpattern $Y\preceq X$, there can be no set of steps which may sort $X$. For every foot-sortable sock ordering $S$, $Y\preceq S$ implies that $Y$ is foot-sortable.
Definition 9.
At a particular sorting stage $(S,R)$, some color $a\in SR$ is sortable if both of the following conditions hold:
  1. (1)
    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top element on the stack.)
  2. (2)
    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct. (In other words, the color $a$ is not blocked by a sandwich.)
Otherwise, $a$ is unsortable in $(S,R)$. We say that a color $a\in X$ is sortable in $X$ if $a$ is sortable in $(\emptyset,X)$. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal.
The following observation follows from the definitions.
Observation 10.
Let $S$ and $R$ be sock orderings. Suppose $x\in SR$ and that $xy\not\subseteq S$ for any $y\in S$ distinct from $x$. Then the color $x$ is sortable in $(S,R)$ if and only if $x$ is sortable in $SR$.
Definition 11.
A color $a$ is a good sortable color at a stage $(S,R)$ if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable.
At times, we may reach a stage in which multiple colors are sortable. For some colors $\{a_{1},\dots,a_{n}\}$ in a string $X$, we say $a_{1}$ is the first terminating color from the set if the final sock of color $a_{1}$ appears before the final sock of color $a_{i}$ for all $a_{i}$ in the set distinct from $a_{1}$.
Definition 12.
A trivially unsortable sock ordering is a sock ordering in which no color is sortable. That is, a sock ordering $X$ is trivially unsortable if for every color $a\in X$, there exist distinct colors $b,c\in X$, different from $a$, such that $bcba\subseteq X$.
We often discuss particular sorting choices in an attempt to sort a sock ordering $X$. A move in some sorting procedure consists of sorting exactly one color. That is, when $(S,R)$ appears, a color $x$, identified as a sortable in $(S,R)$, is chosen to be sorted. Any $x$ are removed from the top of $S$, and if $x\in R$, then for every sock in $R$, we apply the following rule:
  • If the sock is the final sock of color $x$, place the sock into the stack, then remove it from the stack and place the sock on the sorted left side. Immediately terminate the procedure at this point.
  • If the sock has color $x$ but is not the final sock of color $x$, stack the sock and immediately unstack it, bringing it to the sorted left side. Look to the next sock on the right side.
  • If the sock has a color which is not $x$, stack the sock, and do not unstack it. Look to the next sock on the right side.
In contrast, a step in the sorting procedure refers to the movement of exactly one sock: either stacking the first sock appearing on the right, or removing the top element of the stack and bringing it to the sorted left side. In general, it takes multiple steps to complete one sorting move. When we say that we sort the color $a$ we mean that we apply a move as describe above of the sortable color $a$.
Definition 13.
Let $X$ be a sock ordering, and let $\mathcal{A}_{X}$ be the set of distinct colors appearing in $X$. A sorting sequence $\theta$ for $X$ is a nonempty sequence $(a_{1},a_{2},\dots,a_{n})$ of elements from $\mathcal{A}_{X}$, each appearing at most once, such that $a_{1}$ is sortable in $(\emptyset,X)$, and for $i>1$, $a_{i}$ is sortable after $a_{i-1}$ has been sorted. We write $\theta(X)=(S,R)$ if, beginning with $(\emptyset,X)$, $(S,R)$ appears after sorting each color from $\theta$ in sequence.
Observation 14.
At some stage $(S,R)$, if a color $x$ only appears as the top color in $S$, then $x$ is sortable with respect to $(S,R)$. Furthermore, if we sort $x$ to form $(S^{\prime},R)$, then $(S,R)$ is sortable if and only if $(S^{\prime},R)$ is sortable.
Definition 15.
We say that a sorting sequence $\theta$ fully sorts $X$ if $\theta(X)=(\emptyset,\emptyset)$. In light of Observation 14, we also say that $\theta$ fully sorts $X$ if we can fully sort $X$ using the prescribed sorting order $\theta$ with the additional stipulation that we immediately sort a color if it only appears as the top color on the stack. Note that, with this definition, a sorting sequence $\theta$ can fully sort a sequence $X$ even if $\theta$ does not contain all of the colors of $X$.
While a sorting sequence $\theta$ must obey the rules of foot-sorting (that is, $\theta$ may never make a move that causes a sandwich to appear in the stack), there is no requirement for $\theta$ to fully sort $X$.
Example 16.
Taking $\theta_{1}=(a,c,b)$, we have that $\theta_{1}(abcabc)=(\emptyset,\emptyset)$, and thus $\theta_{1}$ fully sorts $abcabc$. Choosing instead $\theta_{2}=(a)$, we have $\theta_{2}(abcabc)=(bc,bc)$.
2.1. The basis of foot-sortable sock orderings
Definition 17.
A minimally unsortable, or minimal, sock ordering $X$ is an unsortable sock ordering in which for all $Y\preceq X$, $Y$ is unsortable $\implies$ $|Y|=|X|$.
Observation 18.
Let $\Gamma$ be the set of all minimally unsortable sock orderings, and let $M$ be the basis of foot-sortable sock orderings, the smallest set of sock orderings which foot-sortable sock orderings must avoid. Then $M=\Gamma$.
As mentioned in the introduction, the basis of foot-sortable sock orderings is infinite. The following definition is used to describe our infinite classes in the basis $\Gamma$.
Definition 19.
Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ distinct colors. Define $I(t_{0}):=I(T)$ to be the empty string when $n=0$, and when $n\geq 1$, let
@@ 2.1. The basis of foot-sortable sock orderings
and it is not hard to see that both statements are true.
@@
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that, when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
@@
leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$ and the conclusion again follows by the induction hypothesis.
@@
where at this one of the two is a strict subordering. By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. First assume that $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$, so $S^{\prime}\neq abt_{i}$. If $b\in S^{\prime}$ and $t_{i}\notin S^{\prime}$, then we sort $b$ and then $t_{i}$. Otherwise, we only sort $t_{i}$. In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a strict subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$. The conclusion then follows from the induction hypothesis.
Now, assume $S^{\prime}=abt_{i}$. So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$ to reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$. In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis. Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a strict subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$. Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis. ∎
We conclude this section by enumerating $\Gamma$, the basis of foot-sortable sock orderings. The following list consists of every sock ordering which is minimally unsortable. We describe five classes of minimally unsortable sock orderings which have infinitely many members, and we also outline other classes of sock orderings which are finite in number. We have divided the basis into categories that correspond with particular subpatterns we will later generate using a greedy sorting algorithm. Sock orderings in classes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have color labels which denote the behavior of particular socks during the greedy sorting algorithm’s procedure.
We have the following orderings in the class $\mathcal{T}$, separated by the number of distinct colors in the sock ordering:
@@
Note that all of the sock orderings in the following class, $\mathcal{C}_{1}$, contain either $atdarat$, $atdrart$, or $atdarbrt$. Furthermore, sorting the color $d$ first for each of these three subpatterns yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and none of these three states are sortable. To then see that the orderings in $\mathcal{C}_{1}$ are unsortable, it is only necessary to verify that in each of these orderings, we must sort $d$ before the colors $a$, $b$, $r$, and $t$.
@@
The other finite class is $\mathcal{C}_{2}$, consisting of those sock orderings which contain either the subordering $abtdabt$ or $abtdatb$. Since $(abt,at)$ is unsortable and $d$ must be sorted before the colors $a$, $b$ and $t$ in the following orderings, it is not difficult to verify that the the following orderings are minimally unsortable.

### v2

@@ Abstract.
Defant and Kravitz considered the following problem: Suppose that, to the right of a foot, there is a line of colored socks that needs to be sorted. However, at any point in time, one can only either place the leftmost sock to the right of the foot onto the foot (stack) or remove the outermost sock on the foot and make it the rightmost sock to the left of the foot (unstack). In this paper, we explicitly describe all minimal initial sock orderings that are unsortable.
TM: Department of Mathematics and Statistics, University of South Florida, Tampa, FL. Email: molla@usf.edu. Research supported in part by NSF grants DMS-1800761 and DMS-2154313.
CN: Department of Mathematics, University of Tennessee, Knoxville, TN. Email: cnelso74@vols.utk.edu. The majority of this research effort was performed while this author was an undergraduate student at the University of South Florida.
@@ 1. Introduction
Consider the following problem, which was introduced by Defant and Kravitz [1]. Starting with an arbitrary line of colored socks to the right of a foot, one aims to line the socks up by color to the left of the foot, but, at any point in time, only one of the following two actions can be performed.
  1. (stack)
    Take the leftmost sock on the right and place it on the foot (potentially over any socks that are already on the foot).
  2. (unstack)
    Remove the outermost sock from the foot and place it so that it becomes the rightmost sock on the left.
The initial ordering of socks on the right is a sock ordering and a sock ordering is foot-sortable (or simply sortable) if, using only the two specified operations, one can produce a sorted ordering of the socks on the left. Otherwise, it is unsortable. Throughout, we represent sock orderings as strings in the obvious way; left-to-right in the string corresponds to left-to-right in the sock ordering. For example, if $r$, $g$, and $b$ refer to socks of color red, green, and blue, respectively, the sock ordering $rgbrgb$ is sortable (to $rrbbgg$ on the left) while the sock ordering $rgbgrgb$ is unsortable.
Let $X$ and $Y$ be sock orderings. We say that $Y$ is a subordering of $X$ and write $Y\subseteq X$ (or $X\supseteq Y$) if deleting zero or more socks from $X$ yields $Y$ exactly. The subordering $Y\subseteq X$ is proper if $X\neq Y$. An unsortable sock ordering is minimally unsortable if every proper subordering is sortable. We write $a\in X$ if the color $a$ appears in $X$ and we let $X-a$ be the subordering of $X$ obtained by deleting all socks of color $a$ from $X$. We use $\emptyset$ to denote the empty ordering and $a$ to denote the ordering consisting of a single sock of color $a$. We let $XY$ be the concatenation of the orderings $X$ and $Y$. We say $X$ and $Y$ are equivalent and write $X\cong Y$ if $Y$ can be obtained from $X$ via a bijective mapping of the colors of $X$ to the colors of $Y$. We call $Y$ a subpattern of $X$ and write $Y\preceq X$ (or $X\succeq Y$) if there exists a sock ordering $Z$ such that $Z\subseteq X$ and $Z\cong Y$. Otherwise, we say that $X$ avoids $Y$. Furthermore, $X$ avoids a collection of sock orderings if it avoids every ordering in the collection.
Modulo the relation $\cong$ and with respect to the partial ordering $\preceq$, Defant and Kravitz called the set of minimal elements among the collection of unsortable sock orderings the basis of foot-sortable sock orderings [1]. One may equivalently view the basis of foot-sortable sock orderings as a collection of pairwise inequivalent minimally unsortable sock orderings that no unsortable sock ordering avoids. In this paper, we give an explicit description of the basis of foot-sortable sock orderings (Section 2.2) and a proof that no unsortable sock ordering avoids this basis (Theorem 3).
1.1. Prior Work
As observed by Defant and Kravitz [1], this problem fits into the rich line of research dealing with sorting procedures in which the employed data structures are restricted. For example, sorting permutations with a stack was considered by Knuth [3]. Defant and Kravitz also considered several variants of this problem, including employing multiple stacks (in particular two stacks or two feet) and restricting the number of socks of a given color. Of particular interest with respect to socks, is the case when there are at most two socks of every color. Independent of our work, Yu explicitly described the basis for foot-sortable orderings in which each color appears at most twice [6]. In doing so, Yu determined that this basis is infinite, and this implies that the basis of foot-sortable sock orderings is also infinite. This answered a question of Defant and Kravitz which our work also addresses. Also independent of our work, Yu devised a fast deterministic algorithm that decides if a sock ordering of length $N$ is sortable in time $O(N\log N)$. Provided that the ordering is sortable, this algorithm can also procedure a valid sorting. This problem was also investigated by Xia, who considered a deterministic algorithm that forces the stack to remain sorted [5].
2. Results
2.1. Preliminaries
Throughout, we will refer to the sorting process using primarily stack-centered language. We denote a particular state of the foot-sorting process as $(S,R)$. Here $S$ represents the contents of the stack, with left-to-right in $S$ corresponding with bottom-to-top in the stack (or inside-to-outside on the foot), while $R$ represents what remains on the right. We say that the state $(S,R)$ is foot-sortable (or just sortable) if we can apply the operations stack and unstack so that all of the socks in $SR$ are arranged by color to the left of the foot. Note that if the stack $S$ is not sorted, then $(S,R)$ is unsortable. As such, we implicitly assume throughout that the stack $S$ is sorted.
Instead of dealing with the two operations stack and unstack, we work almost exclusively at a slightly higher level. At each point, we are mainly concerned with the next sock color that will be moved completely to the left and we say that a color is sortable if it is possible to do this while keeping the stack is sorted. More explicitly, given a state $(S,R)$, a color $a\in SR$ is sortable if the following two conditions hold.
@@ 2.1. Preliminaries
    For any $b\in S$ which is distinct from $a$, $ab\not\subseteq S$. (That is, if $a$ appears on the stack, it must be the top color on the stack.)
@@
    If $a$ appears in $R$, then $bcba\not\subseteq SR$ for any colors $b,c\in SR$ such that $a$, $b$, and $c$ are distinct.
If a color $a\in R$ does not meet the second condition, we say that the color $a$ is blocked by a sandwich. We refer to a sandwich of the form $bcb$ as either a $b$-sandwich or a $c$-sandwich. Given a state $(S,R)$, a color $a\in SR$ is a good sortable color if $a$ is sortable in $(S,R)$ and what remains after sorting $a$ in $(S,R)$ is foot-sortable. If no color is sortable in $(S,R)$, then we say that $(S,R)$ is terminal. Given a sock ordering $X$, a color $a\in X$ is sortable (resp. unsortable) if it is sortable (resp. unsortable) in $(\emptyset,X)$, and a sock ordering $X$ is trivially unsortable if no color in $X$ is sortable. In other words, $X$ is trivially unsortable if the state $(\emptyset,X)$ is terminal.
Note that sorting the color $a$ transitions the state $(S,R)$ to the state $(S^{\prime}R^{\prime},R^{\prime\prime})$ where $S^{\prime}=S-a$ (so $S^{\prime}=S$ if $a$ is not on the stack $S$); and, when $a\in R$, $R^{\prime}$ is the string which precedes the final $a$ in $R$ minus the color $a$ and $R^{\prime\prime}$ is the string which follows the final $a$ in $R$. When $a$ does not appear in $R$, we have $R^{\prime}=\emptyset$ and $R^{\prime\prime}=R$.
A sequence of distinct colors $\theta=a_{1},\dotsc,a_{\ell}$ is a sorting sequence for $(S,R)$ if there are orderings $S_{1},\dotsc,S_{\ell}$ and $R_{1},\dotsc,R_{\ell}$ such that when $S_{0}:=S$ and $R_{0}:=R$, for every $i\in[\ell]$, the color $a_{i}$ is sortable with respect to the state $(S_{i-1},R_{i-1})$ and sorting $a_{i}$ in this state leads to $(S_{i},R_{i})$. We define $\theta(S,R):=(S_{\ell},R_{\ell})$ and we say that $\theta$ fully sorts $X$ if $S_{\ell}=R_{\ell}=\emptyset$. For a sock ordering $X$, we let $\theta(X):=\theta(\emptyset,X)$.
2.2. The basis of foot-sortable sock orderings
We use $\Gamma$ to denote the basis of foot-sortable sock orderings. As mentioned previously, the set $\Gamma$ is infinite and we use the following definition to describe the infinite classes.
@@ 2.2. The basis of foot-sortable sock orderings
Let $T:=(t_{0},\dotsc,t_{n})$ be a sequence of $n+1$ colors. Define $I(t_{0}):=I(T)$ to be the empty ordering when $n=0$, and when $n\geq 1$, let
@@
and it is not hard to see $(abt_{n},at_{n})$ is terminal and removing any sock from $(abt_{n},at_{n})$ makes the state sortable. Therefore, both of the statements are true when $i=n$.
@@
leads to $(abt_{i+1},I(t_{i+1},\dotsc,t_{n})at_{n})$. Note that when $n-i=1$ the only sortable color is $t_{i}$, so the first statement follows by induction. When $n-i\geq 2$, the only sortable colors are $t_{i}$ and $t_{i+1}$ and sorting $t_{i+1}$ in
@@
leads to $(abt_{i}t_{i}t_{i+2},I(t_{i+2},\dotsc,t_{n})at_{n})$. Therefore, the first statement again follows by the induction hypothesis.
@@
where at least one of the two, $S^{\prime}$ or $R^{\prime}$, is a proper subordering. By possibly adding socks, we also assume that either $S^{\prime}=abt_{i}$ or $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. First assume that $R^{\prime}=t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$, so $S^{\prime}\neq abt_{i}$. If $b\in S^{\prime}$ and $t_{i}\notin S^{\prime}$, then we sort $b$ and then $t_{i}$. Otherwise, we only sort $t_{i}$. In either case, we reach $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}$ is a proper subordering of $abt_{t+1}$ and $R^{\prime\prime}=I(t_{i+1},\dotsc,t_{n})at_{n}$. The conclusion then follows from the induction hypothesis.
Now, assume $S^{\prime}=abt_{i}$. So, $R^{\prime}\neq t_{i+1}t_{i}I(t_{i+1},\dotsc,t_{n})at_{n}$. If $R^{\prime}$ does not begin with $t_{i+1}t_{i}$, then we can sort $t_{i}$ and then $b$. After possibly pushing $t_{i+1}$ onto the stack, we arrive at the state $(S^{\prime\prime},R^{\prime\prime})$ where $S^{\prime\prime}\subseteq at_{i+1}$ and $R^{\prime\prime}\subseteq I(t_{i+1},\dotsc,t_{n})at_{n}$. In this case, since $b\notin S^{\prime\prime}$, the conclusion follows from the induction hypothesis. Otherwise, we have $S^{\prime}=abt_{i}$ and $R^{\prime}=t_{i+1}t_{i}R^{\prime\prime}$ where $R^{\prime\prime}$ is a proper subordering of $I(t_{i+1},\dotsc,t_{n})at_{n}$. Then, after sorting the color $t_{i}$, we have $(abt_{i+1},R^{\prime\prime})$ and the conclusion follows by the induction hypothesis. ∎
The following tables describe the basis $\Gamma$. The first table consists of trivially unsortable orderings. It is not difficult to verify that the orderings are minimally unsortable.
@@
Note that all of the sock orderings in the next table have either $atdarat$, $atdrart$, or $atdarbrt$ as a subordering. Furthermore, sorting the color $d$ (preceded or followed by sorting the color $e$ when it appears) for each of these three possible suborderings yields $(at,arat)$, $(at,rart)$, and $(at,arbrt)$, respectively, and each of these three states are terminal. Further note that in the orderings below, $d$ and $e$ are the only sortable colors, which implies that the orderings are unsortable. It is then not difficult to verify that the orderings are minimally unsortable.
@@
The next table consists of the sock orderings that have either $abtdabt$ or $abtdatb$ as a subordering. In a similar fashion to the other classes, sorting $d$ (and possibly $e$) first in these orderings yields some $(S,R)$ where $abt\subseteq S$ and $at\subseteq R)$. Since $(abt,at)$ is unsortable and $d$ and $e$ are the only sortable colors in the following orderings, we have that orderings in this class are unsortable, and it is not difficult to verify that these are minimally unsortable.